# NICER Spectral and Timing Analysis of 4U 1630–47 and its Heartbeat State

### v1

@@ Abstract
We present a spectral and timing analysis of NICER observations of the black hole X-ray binary 4U 1630–47 from 2018 to 2024. We find relativistic reflection features in the hard and soft intermediate states, and disk wind absorption features in the soft intermediate state and soft state. We fit the reflection features with relxillCP and find a stable and untruncated disk in the intermediate states; we fit the wind features with zxipcf and find a stable, highly ionized wind with high column density across different outbursts. Specifically, the heartbeat state is seen in two observations in 2021 and 2023 respectively. Through the phase-resolved spectral fitting, we find the flux to be correlated with the disk parameters while no strong correlation with the coronal parameters is observed, consistent with the scenario given by the inner disk radiation instability. A significant hard lag on the time scale of a second and high coherence is observed near the characteristic frequency of the heartbeat, which can be explained by the viscous propagation of mass accretion fluctuations in the disk. The positive relationship between the heartbeat fractional rms and energy can possibly be explained by a disk-originated oscillation which is then magnified by the corona scattering.
@@ 1 Introduction
Black hole X-ray binaries (BHXRBs) are binary systems in which a central stellar-mass black hole accretes mass from a companion star. During an outburst, BHXRBs are usually divided into different states: the hard state, hard intermediate state (HIMS), soft intermediate state (SIMS) and soft state based on their spectral and timing properties (Homan & Belloni, 2005). The hardness - intensity diagram (HID, Belloni et al., 2011), power color diagram (Heil et al., 2015) and hardness - rms [^1] Referring to the fractional “root mean squared” (rms) noise. diagram (HRD, Plant et al., 2014) are widely used in tracing the distinctive states of BHXRBs. Physically, different states are closely related to changes in the accretion process, as well as the outflows of the system in the form of winds and jets. The dynamic nature of the accretion process makes it possible to study the evolution of accretion geometry and the driving physics behind it (Neilsen et al., 2012).
@@
The underlying mechanism behind the heartbeat is not fully understood. Commonly, the heartbeat mechanism is thought to be associated with a radiation instability of the inner disk (Lightman & Eardley, 1974; Done et al., 2007; Neilsen et al., 2012), a hypothesis which is broadly supported by phase-resolved and time-resolved spectral fitting (Neilsen et al., 2011; Rawat et al., 2022; Belloni et al., 1997). Moreover, the time scale of the modulation is found to be commensurate with the viscous timescale of the inner disk region (Belloni et al., 1997). Wang et al. (2024) further discussed the possibility of the heartbeat arising from the vertical disk instability in systems with a long orbital period (Życki et al., 1999) and disk tearing due to spin-orbit misalignment (Raj & Nixon, 2021) in the case of IGR J17091–3624.
4U 1630–47 is a recurrent transient which typically outbursts every few hundred days (Choudhury et al., 2015; Jones et al., 1976). Previous spectral and timing studies on 4U 1630–47 have shown interesting features of this source: It has a high hydrogen column density along the line of sight, estimated to be of the order $N_{\rm H}\approx 10^{23}\,\mathrm{cm^{-2}}$ with high resolution Chandra data (Gatuzz et al., 2019). Its dimensionless spin parameter has been estimated to be higher than 0.9 through both continuum fitting and reflection modeling (Kushwaha et al., 2023; King et al., 2014; Pahari et al., 2018). The inclination angle is measured to be quite high, around $\sim 64^{\circ}$ through reflection modeling (King et al., 2014). Prominent absorption features from disk winds are commonly seen for 4U 1630–47, e.g., in Pahari et al. (2018); Trueba et al. (2019); Kushwaha et al. (2023); Ratheesh et al. (2024). The source is mainly observed in its soft state. Through spectral and timing analysis of INTEGRAL and RXTE data, Capitanio et al. (2015) find that the hard state is missing in the 2006 and 2008 outbursts. The heartbeat state of this source was first observed in 1998 with RXTE (Trudolyubov et al., 2001). Yang et al. (2022) observe a clear transition from a hard-state with a low-frequency type-C QPO to the (lower-frequency) heartbeat state in 2021 with Insight-HXMT. Notably, for the two heartbeat states of 4U 1630–47 observed previously, the level of luminosity and hardness is similar (Yang et al., 2022).
@@ 2 Observation and Data Reduction
We analyze 251 NICER observations of the source from 2018 to 2024. We perform the standard calibration and screening on the archival NICER data with nicerl2222https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/nicerl2/ in HEASOFT v6.31.1. We use the 3C50 background model (Remillard et al., 2022). We then use the nicerl3-spect [^3] https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/nicerl3-spect/ task to generate the source and background spectra, and the corresponding response files for each observation. The public software stingray [^4] https://docs.stingray.science/en/stable/index.html v2.1 is used to generate the light curve in the 0.2–12 keV energy band (Fig. 1), the hardness-intensity diagram (HID, Fig. 2), and the PDS of the light curves from the original events files. The count rates of different detectors are normalized to 52-FPM equivalent intensity, and we adopt this normalization throughout this work. When computing the PDS, the time resolution of the light curves is 0.005 s, corresponding to a Nyquist frequency of 100 Hz. To create an averaged PDS from different segments of a light curve, the size of each segment is chosen to be 256 s, corresponding to a minimum frequency of about 0.003 Hz, which is enough for the study of heartbeat features at several mHz.
From the light curve, we can see that in the observations in most years, as a new outburst onsets, the count rate first rises over a period of a few weeks and later decreases over a timeframe which can be similar or much longer. Since the HID of this source does not follow the “Q”-shaped track typical of most BHXRB transients, and additionally complicated by the very large column-density in the line of sight, it is challenging to distinguish different states from the HID in isolation. Therefore, we additionally explore the timing-variability features as well as the spectral features, and we show the hardness-rms (see definition in van der Klis, 1995) diagram (HRD) in Fig. 3. Based on classifications of states discussed in Plant et al. (2014) for GX 339–4 with RXTE, Alabarta et al. (2020) for MAXI J1727–203 observed by NICER, Stiele & Kong (2020) for MAXI J1820 observed by NICER and Swift, and more general discussions in Belloni (2010); Belloni & Motta (2016), we show the distinction of different states based on the level of rms on the HRD. We find most of the observations in this 6 year-interval are in the SIMS or soft state. However, in 2021, we detect a clear transition from the HIMS to SIMS.
@@
Figure 4: The PDS of the two observations when the heartbeat is observed. Observation ID 4130010115 corresponds to the observation on September 27, 2021 (MJD 59485), while Observation ID 6130010109 corresponds to the observation on March 27, 2023 (MJD 60031). We can see a clear broad peak at around 0.05 mHz in each, corresponding to a modulation period of about 20 s.
Figure 5: The 300 s light curve segments of the heartbeat states corresponding to the PDS in Fig. 4. The upper panel is the 2021 heartbeat state, and the lower panel shows the 2023 heartbeat state.
@@ 3.1 Simple continuum model
We first fit the spectra with a simple model: tbabs $\times$ diskbb. We employ the diskbb model (Mitsuda et al., 1984) to fit the multi-temperature blackbody component from the accretion disk, the tbabs model to account for interstellar absorption (Wilms et al., 2000) with the column density $N_{\rm H}$ as a free parameter in the fits. We see $N_{\rm H}\approx 10^{23}\,\mathrm{cm^{-2}}$, consistent with earlier studies (Gatuzz et al., 2019; Pahari et al., 2018). While a significant fraction of the observations are successfully fitted with this simple model (i.e., with a reduced $\chi^{2}/\nu\sim 1$),est of the fits yield unacceptable fits with $\chi^{2}/\nu$ even reaching 3-4 in some cases. This indicates that most of the spectra of this source cannot be described well by this simple model, therefore we need to adapt more complex models and take other spectral components like the Comptonization component from the corona into consideration.
@@
Figure 8: A typical and representative spectrum with disk wind features, taken from Observation ID 5130010101 obtained on July 30, 2022 (MJD 59786). The top panel shows the fits with (blue) and without (red) having included a wind component. The two magenta dashed vertical lines mark the 6.69 keV Fe XXV line and 6.97 keV Fe XXVI line. The bottom two panels show the residuals from both fits.
@@
To determine the hydrogen column density and elemental abundances in tbfeo, which should be stable over time, we select nine observations spanning different accretion rates and states (three with low count rate, three intermediate, and three with high count rate. These are indicated by the green triangles in Fig. 1). We fit these simultaneously with the column density and elemental abundances in tbfeo tied together while the disk and coronal parameters are left free for each observation. The joint fit (shown in Fig. 6) gives a reduced $\chi^{2}$ of 743/937 = 0.8 (perhaps indicating that the systematic uncertainties adopted for the NICER processing are overestimated). From this, we obtain a column density of $N_{\mathrm{H}}=16.96^{+0.58}_{-0.41}\times 10^{22}~{}\mathrm{cm^{-2}}$, oxygen and iron abundances of $[O]=0.21^{+0.05}_{-0.06}$ and $[Fe]=0.69^{+0.07}_{-0.05}$, respectively. This fit is shown in Fig 6, from which it can be seen that the model tracks our reference spectra without any obvious outlier observations performing poorly. In our subsequent analysis, we fix the parameters of tbfeo at the values obtained here. However, we note that only fitting data above 2 keV limits our ability to robustly constrain individual elemental abundances. One can reasonably expect that O abundance fitted acts as a proxy for the aggregate of low-Z metals.
@@ 3.2 Disk wind absorption features
We use the multiplicative absorption model zxipcf (Reeves et al., 2008) to model absorption by an ionized wind. There are around 70 spectra in total showing wind absorption features, and they are mostly in the SIMS and soft state (see Fig. 3). The model we adopted to fit these spectra is: tbfeo $\times$ zxipcf $\times$ thcomp $\otimes$ diskbb, and the fitting results are shown in Fig. 10. The covering fraction of the wind over the source is high and touches the upper bound of 1 in most cases (annotated with red arrows). The column density of the wind is ${\rm log}\,N_{\rm H,wind}\approx 24\pm 1\,\mathrm{cm^{-2}}$, and the ionization parameter $\log\xi\approx 5\pm 1$. From year-to-year, no significant evolution is apparent in the properties of the wind. Our results are consistent with those obtained using a different model by Ratheesh et al. (2024), in which they use the absorption model cloudy (Ferland et al., 2017) and also find that $N_{\mathrm{H}}\approx 10^{24}\,\mathrm{cm^{-2}}$ and $\xi\approx 10^{5}$.
Figure 10: Fitting results of the parameters in the model tbfeo $\times$ zxipcf $\times$ thcomp $\otimes$ diskbb. The parameters are (from top to bottom): the covering fraction, the column density, and the ionization degree of the wind. Points touching the upper bound of the fit are annotated with red arrows.
@@ 3.3 Relativistic reflection features
We use the relativistic reflection model relxillCp (García et al., 2014) to fit the the reflection features. From 2018 to 2024, there are only 9 NICER observations showing a clear reflection component, which are observed during the HIMS to SIMS state transition at the beginning of the 2021 outburst. The model used to fit these spectra is tbfeo $\times$ thcomp $\otimes$ (diskbb + relxillCp). To avoid the degeneracy between the inner disk radius and the spin (Buisson et al., 2019), we fix the spin at the maximum value 0.998 in relxillCp and study the change of $R_{\mathrm{in}}$. The inclination angle is fixed at $64^{\circ}$ (King et al., 2014). The reflection fraction is set to $-1$ so that only the reflection component is produced (noting that the Comptonized emission is already described by thcomp $\otimes$ diskbb). The coronal temperature $kT_{\mathrm{e}}$ and photon index $\Gamma$ in the Comptonization component and reflection component are linked together. Due to the lack of data higher than 10 keV, the corona temperature cannot be constrained well from the fits, so we fix it at a reasonable benchmark of 50 keV. The emissivity of the reflection component is modeled with a power law ($\epsilon\propto r^{-q}$).
@@ 3.4 Fitting results of the thermal component
We summarize the results of fitting the thermal component with diskbb for all the observations in Fig. 12. Generally, the inner disk temperature is anti-correlated with the diskbb normalization factor. At the end of the 2020 observations and the 2019 observations, the inner disk temperature decreases to a quite low value at around 0.2 keV and the normalization factor of diskbb increases when the source returns from the soft state to the hard state (see lower right points in the HID in Fig. 2). The estimates for these lower-temperature points ($T_{\rm{in}}\lesssim$ 0.4 keV) have large systematic uncertainties due to the exclusion of low-energy data and the low total counts in these spectra. When the source becomes brighter in the soft state, the inner disk temperature and the diskbb normalization factor change in an inverse manner. For most of the time, the inner disk temperature is higher than 1 keV, indicating the dominance of the soft state of the source, which is also found in Capitanio et al. (2015) through spectral-timing analysis of the 2006, 2008 and 2010 outbursts of 4U 1630–47.
It is worth noting that despite the similarity in the hardness, luminosity and timing properties between the two heartbeat states, we find they have quite different spectra and thermal features. We note that for both the 2021 and 2023 outburst, sudden and large-amplitude changes in the inner disk temperature before, during and after the heartbeat state are observed (further details about the spectral and temperature change can be found in Appendix B). In Section. 6, we explore possible explanations for this large temperature shift.
Figure 12: The fitting results showing the relationship between the inner-disk temperature and the disk normalization. The dashed line show constant $R_{\mathrm{in}}$ contours, calculated from the equation $norm_{\mathrm{diskbb}}=(R_{\mathrm{in}}/f_{\mathrm{col}}^{2}D_{10})^{2}\cos{i}$ (Kubota et al., 1998) and adopting a temperature-dependent color correction factor $f_{\mathrm{col}}\propto T_{\mathrm{in}}^{1/4}$ (Davis et al., 2006).

### v2

@@ Abstract
We present a spectral and timing analysis of NICER observations of the black hole X-ray binary 4U 1630–47 from 2018 to 2024. We find relativistic reflection features in the hard and soft intermediate states, and disk wind absorption features in the soft intermediate state and soft state. We fit the reflection features with relxillCP and find a stable and untruncated disk in the intermediate states; we fit the wind features with XSTAR and find a stable, highly ionized wind with high column density across different outbursts. Specifically, the heartbeat state is seen in two observations in 2021 and 2023 respectively. Through the phase-resolved spectral fitting, we find the flux of the source to be correlated with the disk parameters while no strong correlation with the coronal parameters is observed, consistent with the scenario given by the inner disk radiation pressure instability. A hard lag on the time scale of $1~{}\rm{s}$ and high coherence is observed near the characteristic frequency of the heartbeat, which can be explained by the viscous propagation of mass accretion fluctuations in the disk. The positive relationship between the heartbeat fractional rms and energy can possibly be explained by a disk-originated oscillation which is then magnified by the corona scattering.
@@ 1 Introduction
Black hole X-ray binaries (BHXRBs) are binary systems in which a central stellar-mass black hole accretes mass from a companion star. During an outburst, BHXRBs are usually divided into different states: the hard state, hard intermediate state (HIMS), soft intermediate state (SIMS) and soft state based on their spectral and timing properties (Homan & Belloni, 2005). The hardness - intensity diagram (HID, Belloni et al., 2011), power color diagram (Heil et al., 2015) and hardness - rms [^1] Referring to the fractional “root mean squared” (rms) noise. The definition of rms can be found in van der Klis (1995). diagram (HRD, Plant et al., 2014) are widely used in tracing the distinctive states of BHXRBs. Physically, different states are closely related to changes in the accretion process, as well as the outflows of the system in the form of winds and jets. The dynamic nature of the accretion process makes it possible to study the evolution of accretion geometry and the driving physics behind it (Neilsen et al., 2012).
@@
The underlying mechanism behind the heartbeat is not fully understood. Commonly, the heartbeat mechanism is thought to be associated with a radiation pressure instability of the inner disk (Lightman & Eardley, 1974; Done et al., 2007; Neilsen et al., 2012), a hypothesis which is broadly supported by phase-resolved and time-resolved spectral fitting (Neilsen et al., 2011; Rawat et al., 2022; Belloni et al., 1997). Moreover, the time scale of the modulation is found to be commensurate with the viscous timescale of the inner disk region (Belloni et al., 1997). Wang et al. (2024) further discussed the possibility of the heartbeat arising from the vertical disk instability in systems with a long orbital period (Życki et al., 1999) and disk tearing due to spin-orbit misalignment (Raj & Nixon, 2021) in the case of IGR J17091–3624.
4U 1630–47 is a recurrent transient which typically outbursts every few hundred days (Choudhury et al., 2015; Jones et al., 1976). Previous spectral and timing studies on 4U 1630–47 have shown interesting features of this source: It has a high hydrogen column density along the line of sight, estimated to be of the order $N_{\rm H}\approx 10^{23}\,\mathrm{cm^{-2}}$ with high resolution Chandra data (Gatuzz et al., 2019). Its dimensionless spin parameter has been estimated to be higher than 0.9 through both continuum fitting and reflection modeling (Kushwaha et al., 2023; King et al., 2014; Pahari et al., 2018), but see also Liu et al. (2022). The inclination angle is measured to be quite high, around $\sim 64^{\circ}$ through reflection modeling (King et al., 2014). Prominent absorption features from disk winds are commonly seen for 4U 1630–47, e.g., in Pahari et al. (2018); Trueba et al. (2019); Kushwaha et al. (2023); Ratheesh et al. (2024). The source is mainly observed in its soft state. Through spectral and timing analysis of INTEGRAL and RXTE data, Capitanio et al. (2015) find that the hard state is missing in the 2006 and 2008 outbursts. The heartbeat state of this source was first observed in 1998 with RXTE (Trudolyubov et al., 2001). Yang et al. (2022) observe a clear transition from a hard-state with a low-frequency type-C QPO to the (lower-frequency) heartbeat state in 2021 with Insight-HXMT. Notably, for the two heartbeat states of 4U 1630–47 observed previously, the level of luminosity and hardness is similar (Yang et al., 2022).
@@ 2 Observation and Data Reduction
We analyze 251 NICER observations of the source from 2018 to 2024. During these years, the NICER calibration is not known to have changed appreciably. But in May 2023, there was an increase in optical loading on the instrument, and as a result, during the ISS daytime, increased noise can affect the study of low-energy photons below 1 keV [^2] https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/light-leak-overview. However, this is unlikely to influence our study on 4U 1630–47, because only higher energy photons are considered in spectral and timing analysis due to the high interstellar absorption.
We perform the standard calibration and screening on the archival NICER data with nicerl2333https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/nicerl2/ in HEASOFT v6.31.1. We use the 3C50 background model (Remillard et al., 2022). We then use the nicerl3-spect [^4] https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/nicerl3-spect/ task to generate the source and background spectra, and the corresponding response files for each observation. The public software stingray [^5] https://docs.stingray.science/en/stable/index.html v2.1 is used to generate the light curve in the 0.2–12 keV energy band (Fig. 1), the hardness-intensity diagram (HID, Fig. 2), and the PDS of the light curves from the original events files. Throughout the paper, uncertainties are reported at the 90% confidence level. The count rates of different detectors are normalized to 52-FPM equivalent intensity, and we adopt this normalization throughout this work. When computing the PDS, the time resolution of the light curves is 0.005 s, corresponding to a Nyquist frequency of 100 Hz. To create an averaged PDS from different segments of a light curve, the size of each segment is chosen to be 256 s, corresponding to a minimum frequency of about 0.003 Hz, which is enough for the study of heartbeat features at several mHz.
From the light curve, we can see that in the observations in most years, as a new outburst onsets, the count rate first rises over a period of a few weeks and later decreases over a timeframe which can be similar or much longer. Since the HID of this source does not follow the “Q”-shaped track typical of most BHXRB transients, and additionally complicated by the very large column-density in the line of sight, it is challenging to distinguish different states from the HID in isolation. Therefore, we additionally explore the timing-variability features as well as the spectral features, and we show the hardness-rms diagram (HRD) in Fig. 3. Based on classifications of states discussed in Plant et al. (2014) for GX 339–4 with RXTE, Alabarta et al. (2020) for MAXI J1727–203 observed by NICER, Stiele & Kong (2020) for MAXI J1820 observed by NICER and Swift, and more general discussions in Belloni (2010); Belloni & Motta (2016), we show the distinction of different states based on the level of rms on the HRD. We find most of the observations in this 6 year-interval are in the SIMS or soft state. However, in 2021, we detect a clear transition from the HIMS to SIMS.
@@
Figure 4: The PDS of the two observations when the heartbeat is observed. Observation ID 4130010115 corresponds to the observation on September 27, 2021 (MJD 59485), while Observation ID 6130010109 corresponds to the observation on March 27, 2023 (MJD 60031). For the 2021 heartbeat, the PDS peaks at $5.6^{+0.4}_{-0.6}$ mHz. For the 2023 heartbeat, the PDS peaks at $4.7^{+0.4}_{-0.4}$ mHz.
Figure 5: The 300 s light curve segments of the heartbeat states corresponding to the PDS in Fig. 4. The upper panel is the 2021 heartbeat state, with an oscillation period of about 18 s (measured from phase-folding). The lower panel shows the 2023 heartbeat state, with an oscillation period of about 20 s (measured from phase folding).
@@ 3.1 Simple continuum model
We first fit the spectra with a simple model: tbabs $\times$ diskbb. We employ the diskbb model (Mitsuda et al., 1984) to fit the multi-temperature blackbody component from the accretion disk and the tbabs model to account for interstellar absorption (Wilms et al., 2000) with the column density $N_{\rm H}$ as a free parameter in the fits. We find $N_{\rm H}\approx 10^{23}\,\mathrm{cm^{-2}}$, consistent with earlier studies (Gatuzz et al., 2019; Pahari et al., 2018). While a significant fraction of the observations are successfully fitted with this simple model (i.e., with a reduced $\chi^{2}/\nu\sim 1$), most of the fits yield unacceptable fits with $\chi^{2}/\nu$ even reaching 3-4 in some cases. This indicates that most of the spectra of this source cannot be described well by this simple model, therefore we need to adapt more complex models and take other spectral components like the Comptonization component from the corona into consideration.
@@
Figure 8: A typical and representative spectrum with disk wind features, taken from Observation ID 5130010101 obtained on July 30, 2022 (MJD 59786). The top panel shows the fits with (blue) and without (red) having included a wind component. The two magenta dashed vertical lines mark the 6.69 keV Fe XXV line and 6.97 keV Fe XXVI line. The two bottom panels show the residuals from both fits.
@@
To determine the hydrogen column density and elemental abundances in tbfeo, which should be stable over time, we select nine observations spanning different accretion rates and states (three with low count rate, three intermediate, and three with high count rate. These are indicated by the green triangles in Fig. 1). We fit these simultaneously with the column density and elemental abundances in tbfeo tied together while the disk and coronal parameters are left free for each observation. The joint fit (shown in Fig. 6) gives a reduced $\chi^{2}$ of 743/937 = 0.8 (perhaps indicating that the systematic uncertainties adopted for the NICER processing are overestimated). From this, we obtain a column density of $N_{\mathrm{H}}=16.96^{+0.58}_{-0.41}\times 10^{22}~{}\mathrm{cm^{-2}}$, and oxygen and iron abundances of $[O]=0.21^{+0.05}_{-0.06}$ and $[Fe]=0.69^{+0.07}_{-0.05}$, respectively. This fit is shown in Fig 6, from which it can be seen that the model tracks our reference spectra without any obvious outlier observations performing poorly. In our subsequent analysis, we fix the parameters of tbfeo at the values obtained here. However, we note that only fitting data above 2 keV limits our ability to robustly constrain individual elemental abundances. One can reasonably expect that O abundance fitted acts as a proxy for the aggregate of low-Z metals.
@@ 3.2 Disk wind absorption features
We calculate a grid model with XSTAR (Kallman & Bautista, 2001) and use the table model to fit the absorption by an ionized wind. The input continuum spectrum of the XSTAR calculation is the averaged best-fit spectrum with the continuum model. Becuase of the similar thermal properties of all the spectra showing wind absorption features (see Fig. 3), almost all spectra can be fit well with the XSTAR table model with the same seed spectrum. The model we adopted to fit these spectra is: tbfeo $\times$ xstar $\times$ thcomp $\otimes$ diskbb, and the three free fitting parameters include the absorber column density ($N_{\rm H,wind}$), the ionization degree ($\log\xi$), and the redshift ($z$). The fitting results are shown in Fig. 10. From year-to-year, no significant evolution is apparent in the properties of the wind. On average, the column density of the wind is $N_{\rm H,wind}\approx(1.9\pm 0.9)\times 10^{23}\,\mathrm{cm^{-2}}$, the ionization parameter $\log\xi\approx 5\pm 2$, and the redshift is $z\approx(-0.003\pm 0.002)$.
Figure 10: Fitting results of the parameters in the model tbfeo $\times$ xstar $\times$ thcomp $\otimes$ diskbb. The parameters are (from top to bottom): the red shift, the column density, and the ionization degree of the wind.
@@ 3.3 Relativistic reflection features
We use the relativistic reflection model relxillCp (García et al., 2014) to fit the the reflection features. From 2018 to 2024, there are only 9 NICER observations showing a clear reflection component, which are observed during the HIMS to SIMS state transition at the beginning of the 2021 outburst. The model used to fit these spectra is tbfeo $\times$ thcomp $\otimes$ (diskbb + relxillCp). To avoid the degeneracy between the inner disk radius and the spin (Buisson et al., 2019), we fix the spin at the maximum value 0.998 in relxillCp and study the change of $R_{\mathrm{in}}$. The inclination angle is fixed at $64^{\circ}$ (King et al., 2014). The reflection fraction is set to $-1$ so that only the reflection component is produced (noting that the Comptonized emission is already described by thcomp $\otimes$ diskbb). The coronal temperature $kT_{\mathrm{e}}$ and photon index $\Gamma$ in the Comptonization component and reflection component are linked together. Due to the lack of data higher than 10 keV, the corona temperature cannot be constrained well from the fits, so we fix it at a reasonable benchmark of 50 keV. We also checked that a different corona temperature does not significantly impact the fit results. For example, when we fix the corona temperature at 100 keV instead, only the absolute values of the photon index $\Gamma$ become slightly lower, while the trends and values of other parameters do not change much. The emissivity of the reflection component is modeled with a power law ($\epsilon\propto r^{-q}$).
@@ 3.4 Fitting results of the thermal component
We summarize the results of fitting the thermal component with diskbb for all observations in Fig. 12. Generally, the inner disk temperature is anti-correlated with the diskbb normalization factor. At the end of the 2020 observations and the 2019 observations, the inner disk temperature decreases to a quite low value at around 0.2 keV and the normalization factor of diskbb increases when the source returns from the soft state to the hard state (see the lower right points in the HID in Fig. 2). The estimates for these lower-temperature points ($T_{\rm{in}}\lesssim$ 0.4 keV) have large systematic uncertainties due to the exclusion of low-energy data and the low total counts in these spectra. When the source becomes brighter in the soft state, the diskbb normalization decreases with increasing temperature. For most of the time, the inner disk temperature is higher than 1 keV, indicating the dominance of the soft state of the source, which is also found in Capitanio et al. (2015) through spectral-timing analysis of the 2006, 2008 and 2010 outbursts of 4U 1630–47.
It is worth noting that despite the similarity in the hardness, luminosity and timing properties between the two heartbeat states, we find they have quite different spectra and thermal features. We note that for both the 2021 and 2023 outbursts, sudden and large-amplitude changes in the inner disk temperature before, during and after the heartbeat state are observed (further details about the spectral and temperature changes can be found in Appendix B). In Section. 6, we explore possible explanations for this large temperature shift.
Figure 12: The fitting results showing the relationship between the inner-disk temperature and the disk normalization. The two arrows show the changing trend of hardness. The dashed line show constant $R_{\mathrm{in}}$ contours, calculated from the equation $norm_{\mathrm{diskbb}}=(R_{\mathrm{in}}/(f_{\mathrm{col}}^{2}D_{10}))^{2}\cos{i}$ (Kubota et al., 1998) and adopting a temperature-dependent color correction factor $f_{\mathrm{col}}\propto T_{\mathrm{in}}^{1/4}$ (Davis et al., 2006).